A boat weighs 1500lbs more then its motor and 1900lbs more then its trailer. Together the boat and motor weigh five times as much as the trailer. How much does the boat weigh?
Answer by nerdybill(7384)

A boat weighs 1500lbs more then its motor and 1900lbs more then its trailer. Together the boat and motor weigh five times as much as the trailer. How much does the boat weigh?
.
Let b = weight of boat
then from: "A boat weighs 1500lbs more then its motor"
b-1500 = weight of motor
and from: "and 1900lbs more then its trailer"
b-1900 = weight of trailer
.
To derive our equation, we use "Together the boat and motor weigh five times as much as the trailer"
b + (b-500) = 5(b-1900)
b + b - 500 = 5b - 9500
2b - 500 = 5b - 9500
-500 = 3b - 9500
9000 = 3b
3000 lbs = b (weight of boat)